All freshmen in a particular school were found to have variation in grades expressed as a standard deviation. Two samples among these freshmen, made up of 30 and 50 students each, were found to have means of 90 and 86 with a standard deviation of 2.75 and 1.50 , respectively. Based on their grades, is the first group really brighter than the second group at .05 level of significance? 


1
Expert's answer
2021-01-20T03:25:21-0500

The following null and alternative hypotheses need to be tested:

"H_0: \mu_1\leq\mu_2"

"H_1: \mu_1>\mu_2"

This corresponds to a right-tailed test, for which a t-test for two population means, with two independent samples, with unknown population standard deviations will be used.

Testing for Equality of Variances

A F-test is used to test for the equality of variances. The following F-ratio is obtained:


"F=\dfrac{s_1^2}{s_2^2}=\dfrac{2.75^2}{1.5^2}=3.361"

"df_1=n_1-1=29, df_2=n_2=1=49, \alpha=0.05"

The critical values are "F_L=0.502" and "F_U=1.881."

Since "F=3.361," then the null hypothesis of equal variances is rejected.

This t-statistic follows a t-distribution approximately, with estimated degrees of freedom


"\dfrac{(\dfrac{s_1^2}{n_1}+\dfrac{s_2^2}{n_2})^2}{\dfrac{1}{n_1-1}(\dfrac{s_1^2}{n_1})^2+\dfrac{1}{n_2-1}(\dfrac{s_2^2}{n_2})^2}"

"=\dfrac{(\dfrac{2.75^2}{30}+\dfrac{1.50^2}{50})^2}{\dfrac{1}{30-1}(\dfrac{2.75^2}{30})^2+\dfrac{1}{50-1}(\dfrac{1.50^2}{50})^2}=39.532,"


Based on the information provided, the significance level is "\alpha=0.05," and the degrees of freedom are "df=39.532," , assuming that the population variances are unequal.

It is found that the critical value for this right-tailed test is "t_c=1.6843," for "\alpha=0.05" and "df=39.532."

The rejection region for this right-tailed test is "R=\{t: t>1.6843\}." 

Since it is assumed that the population variances are equal, the t-statistic is computed as follows:


"t=\dfrac{\bar{X_1}-\bar{X_2}}{\sqrt{\dfrac{s_1^2}{n_1}+\dfrac{s_2^2}{n_2}}}"

"=\dfrac{90-86}{\sqrt{\dfrac{2.75^2}{30}+\dfrac{1.5^2}{50}}}=7.3387"

Since it is observed that "t=7.3387>1.6843=t_c," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that population mean "\mu_1" is greater than "\mu_2," at the 0.05 significance level.


Using the P-value approach: The p-value is "p=0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that population mean "\mu_1" is greater than "\mu_2," at the 0.05 significance level.


Therefore, there is enough evidence to claim that the first group really brighter than the second group, at the 0.05 significance level.